OpenStudy (anonymous):

Please Solve (Find n) 5n+7=2n-3

OpenStudy (anonymous):

n=-10/3

OpenStudy (anonymous):

5n+7=2n-3 5n-2n=-7-3 3n=-10 divide both sides by 3 so.. n=10/3

OpenStudy (anonymous):

that should be negative 10/3

jhonyy9 (jhonyy9):

3n=-10 n=-10/3